daughter needs help, and so do i apparently, sure would like to know the procedure for solving this: 8y+5z=4390
y+z= 680 , any help is more that appreciated,
thanx phil

Answer by stanbon(75887)

8y+5z=4390
y+z = 680
----------------
This pair of equations is called a "system" of equations.
We are looking for a pair of numbers that can replace
y and z and will make both eqations true.
---------------------
One way to "solve" a system like this is called substitution.
The 2nd equation can be written as follows: y = 680-z
------------
Substitute this value of "y" for y in the 1st equation to get:
8(680-z)+5z=4390
5440-8z+5z = 4390
-3z = -1050
z = 350
-----------
Now substitute that value into y = 680-z to solve for "y".
y = 680-350 = 330
-------------------------
Check this solution in the original equations:
8y+5z=4390
y+z= 680
--------
350+330= 680
That's true
----
8*330+5*350 = 4390
That's true
--------
So the solution is y=330 and z=350
===================
Cheers,
Stan H.